OpenStudy (anonymous):

What is the range of the graph of y = 5(x - 2)2 + 7? help?

OpenStudy (anonymous):

y greater than or equal to 7 y less than or greater to 7 y greater than or equal to 2 y less than or greater to 2 answer choices!

OpenStudy (anonymous):

I don't get it:((

OpenStudy (jdoe0001):

look at the "vertex form" of the parabola now look at what you have $$ y = 5(x - \color{red}{2})^2 + \color{red}{7}\\ y = a(x-\color{red}{h})^2 + \color{red}{k} $$ now what do you think is the vertex for your parabola there?

OpenStudy (anonymous):

h?

OpenStudy (jdoe0001):

well the actual values, the (h, k)

OpenStudy (anonymous):

so its 7..but what about the signs?

OpenStudy (jdoe0001):

is at (2, 7) the parabola is going upwards, so it looks like 

OpenStudy (anonymous):

okaaaaaaay...

OpenStudy (jdoe0001):

so what do you think is the range? range meaning the "y" values takes notice, "y" doesn't go below the "vertex"

OpenStudy (anonymous):

is it A?:/

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

AH ! thank you so much!

OpenStudy (jdoe0001):

yw